OpenStudy (anonymous):

MEDAL which expression is equivalent to the one shown below? (4z)^-3 -4z^3 - i think its this -12z^3 1/12z^3 1/64z^3

OpenStudy (anonymous):

\[(4z)^{-3}\]\[\frac{1}{(4z)^{3}}\]\[\frac{1}{4z \times 4x \times 4z}\]\[\frac{1}{64z^{3}}\]